Due to ASF issue it was aired that the average price of a kilo of pork in MM is P195.00. However, a sample of 15 prices randomly collected from different markets in MM showed an average of P200.00 and with standard deviation of P9.50. using 0.05 significant level. What would be your conclusion for the price of pork in MM.



1
Expert's answer
2021-06-07T09:37:36-0400

"t=\frac{\overline{x}-\mu}{\sigma/\sqrt{n}}=\frac{200-195}{9.5/\sqrt{15}}=2.038"

"df=n-1=14"

Critical value:

"t_{.975}=2.145"

Since t < critical value, we can conclude that the claimed average price of a kilo of pork in MM of P195.00 is true.